ASSESSING THE PERFORMANCE OF EMPIRICAL BAYES ESTIMATORS

J. S. MARITZ^{1} AND T. LWIN^{2}

1Department of Statistics, La Trobe University, Bundoora, Victoria 3083, Australia

2Division of Mathematics and Statistics, CSIRO, Private Bag 10, Clayton, Victoria 3168, Australia

(Received June 17, 1991; revised January 27, 1992)

Abstract. Methods for deriving empirical Bayes estimators are generally available. Corresponding general techniques for assessing the performance of these estimators are not widely developed yet, however. In this paper we provide a general procedure for assessing and comparing the performance of the empirical Bayes estimators and other estimators in a given data set.

Key words and phrases: Unequal component sample sizes, empirical Bayes, linear EB, assessment of performance, comparison with MLE.